Life Insurance: Your company sells life insurance. You charge a 55 year old man $70 for a one year, $100,000 policy. If he dies over the course of the next year you pay out $100,000. If he lives, you keep the $70. Based on historical data (relative frequency approximation) the average 55 year old man has a 0.9998 probability of living another year.

(a) What is your expected profit on this policy?
$

(b) What is an accurate interpretation of this value?

It represents the average profit per policy sold that you would expect if you sold a lot of these policies.

It represents the loss on every policy sold.    

It is meaningless because the insurance company never makes this amount on a policy.

It represents the profit on every policy sold
Answer #1

of x be the profit pes policy, P(x = 70)= 0.9998, PCX=- 100000)-0.0002 @ E(X)=$PX; -70*0.9998-100000~0.000 2 = 49.986 $ o Ot

If satisfied with the solution, please rate or do comment in comment section for review or further clarification. Thank you